OpenStudy (anonymous):

Eva created a coffee blend for her restaurant by mixing Kenya and French Roast coffee beans. The Kenya beans cost $15 per pound and the French Roast costs $8 per pound. She bought a total of 10 pounds of coffee for a total of $97.50. How much of each type of bean did she buy? Select one: a. 2.5 lbs Kenya and 7.5 lbs French Roast b. 1.5 lbs Kenya and 8.5 lbs French Roast c. 4 lbs Kenya and 6 lbs French Roast d. 3 lbs Kenya and 7 lbs French Roast

OpenStudy (anonymous):

x=kenya beans pounds y=french roasts pounds x+y=10 15x+8y=97.50 solve system of equations. y=10-x 15x+8(10-x)=97.50 7x+80=97.50 7x=17.5 x=2.5. plug into y=10-x. y=7.5. Answer is A 2.5 lbs Kenya and 7.5 lbs French roast.

OpenStudy (anonymous):

so how did you get 7x

OpenStudy (anonymous):

Closest one to it

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

I got 7.5x not 7

OpenStudy (anonymous):

wait nvm im confused ahah sorry x

OpenStudy (anonymous):

oh thats ok but thnks